Write a liner equation (slope-intercept) perpendicular to 2x+3y=-9 and passing through (-4,5).
Answer by lwsshak3(11628)

Write a liner equation (slope-intercept) perpendicular to 2x+3y=-9 and passing through (-4,5).
***
slope intercept equationof a line: y=mx+b) m=slope, b=y-intercept
For given line:
2x+3y=-9
3y=-2x-9
y=-2x/3-3
m=-2/3
..
For line perpendicular to given line:
slope=3/2(negative reciprocl)
Equation: y=3x/2+b
solve for b using given coordinates(-4,5) on the line
5=3*(-4)/2+b
b=11
equation of perpendicular line: y=3x/2+11